# A Systematic Literature Review of Software Engineering Research on Jupyter Notebook

### v1

@@ 4.4 Industry Collaborations
Software engineering research on Jupyter notebooks is regularly done in collaboration with industry. Of the 146 primary studies on software engineering research on Jupyter notebooks, 31 (21.2%) were done with industry collaborations. Table 3 presents the industry affiliations of the studies included in our SLR. Microsoft Research has the highest representation among the identified industry collaborations, with a study count of 12. IBM Research has six studies on software engineering research on notebooks, followed by Microsoft. We analyzed those studies and found that IBM research mainly focused on code documentation in Jupyter Notebook [91, 101, 174, 169, 172]. On the other hand, the Microsoft Research team focuses on code management [55, 155, 118, 100, 81, 97]. We also found that JetBrains Research was involved in 5 studies, whereas Google and Fujitsu Research each have two studies. Overall, the industry affiliations reflect the widespread interest and impact of software engineering research on Jupyter notebooks in both academia and industry.
@@
Microsoft Research [155, 118, 55, 100, 81, 97, 189, 21, 20, 59, 83, 107]
IBM Research [91, 101, 169, 172, 174, 25]
JetBrains Research [46, 159, 45, 47, 160]
Fujitsu Research of America [10, 203]
Google Inc [114, 201]
Apple [23]
Adobe Research [22]
Megagon Labs [77]
Ploomber [99]
@@ 4.5 Replication Packages
Our systematic literature review discovered a gap in the availability of the resources necessary to replicate the research. Specifically, almost half of the studies (79 out of 146) did not provide publicly accessible source code or installer files of their work. This made it difficult to validate and build on their findings. In contrast, the remaining 67 studies provide source code, installer files, datasets, or a combination thereof to replicate their research results. Upon a closer manual examination of these studies, we found that replication packages for 64 studies are available because they provide the source code with clear instructions to execute. The remaining three studies lacked available replication packages due to the absence of source code or lack of clear execution instructions [146, 131, 145]. For example, Settewong et al. [145] shared a GitHub URL [^10] https://github.com/NAIST-SE/VizJupyterNotebooks of their work; however, upon examination, we found neither source code nor installation files.
Sharing replication packages of software engineering research on notebooks regularly does not follow open science principles. We found that the dominant platform of choice was GitHub, with 54 studies opting to use it to host their replication packages. Although GitHub is a widely accessible platform, any changes or deletions by repository owners can be a risk to the future accessibility of the replication packages. Best practices for open science in software engineering [98] recommend to adopt platforms and repositories designed for the long-term archiving of research outputs. We found that only 13 studies hosted their replication packages on platforms that align with these best practices (i.e., they ensure the permanent archiving of digital artifacts). Of these, 11 were hosted on Zenodo [105, 127, 115, 73, 120, 121, 7, 43, 40, 26, 182] and two on Figshare [90, 147].
@@ 4.6 Jupyter Notebook Extensions
We found that a modest 10.9% of the studies (16 of 146) explored the development and use of Jupyter Notebook extensions to enhance the user experience within notebook environments. Despite showing the potential benefits of Jupyter Notebook extensions, five studies did not provide available replication packages due to missing links or code [25, 101, 186, 168, 81]. We found that eight extensions were reusable and integrable into Jupyter notebooks, showcasing a variety of functionalities designed to improve the user experience of Jupyter notebooks [183, 10, 64, 116, 55, 68, 127, 88]. After replicating and integrating these extensions into Jupyter notebooks, we found that the majority of Jupyter Notebook extensions are designed to enhance the visualization and presentation of notebook cells and their outputs [183, 10, 127, 64, 55, 111]. However, only a few studies tackle more complex challenges related to the core functionalities of Jupyter notebooks. For example, improving the quality and reproducibility of Jupyter notebooks [116], managing data provenance through their extension [68, 37], allowing live migration of notebook environments [88], and notifying potential fairness in data science code [54].
@@
Code reuse and provenance (Section 5.1) Code cloning [73, 198, 63, 131]
Reusing code snippets by code search [56, 84, 186, 85, 10, 57, 7, 126, 83]
Reproducibility [116, 115, 143, 133, 177, 178, 139, 3, 138]
Provenance [74, 117, 28, 68, 69, 137, 66, 37, 54, 154]
Managing computational environment and workflow (Section 5.2) Empirical studies on workflows [78, 124, 24, 118, 130, 75, 156, 43, 205]
Computational environment in notebooks [31, 25, 21, 92, 88, 189, 128, 70, 16, 160, 86, 204, 141, 87]
Managing dependencies [203, 180, 35, 185]
Performance analysis [190, 191]
Readability of notebooks (Section 5.3) Refactoring [90, 140, 29, 30, 159, 146, 55]
Non-linear visualization [183, 64, 134, 20]
@@
Documentation of notebooks (Section 5.4) Empirical studies on documentation [136, 135, 125]
Documentation generation [174, 169, 91, 172, 100, 101, 81, 89, 41, 175, 107]
Cell header generation [167, 165, 166, 111]
@@
Testing and debugging (Section 5.5) Empirical studies on testing and debugging [26, 147, 182]
Detecting bugs [132, 196, 12, 109, 34, 181, 47, 45]
Detecting data leakage [197, 155, 103]
Visualization in notebooks (Section 5.6) Empirical studies on visualizations [2, 145, 127, 184, 194]
Interactive visualization [77, 195, 144, 60, 50]
Best practices (Section 5.7) Following code style standards [46, 179, 149, 15, 122, 123, 152]
Best practices for collaborative use [119, 121, 170]
Cell execution order (Section 5.8) - [105, 151, 61, 99, 93, 14]
AI-based coding assistance for notebooks (Section 5.9) - [201, 13, 97, 22, 38, 58, 23, 188, 76]
Supporting other programming paradigms (Section 5.10) - [104, 114, 168, 187]
Datasets of notebooks (Section 5.11) - [46, 121, 84, 91, 180, 120, 73, 115, 125, 26, 90, 127, 40, 147, 43, 58, 47, 182, 7, 39]
@@ 5.1 Code Reuse and Provenance
Code reuse in Jupyter notebooks refers to the intentional act of leveraging previously written code cells. Code reuse within Jupyter notebooks differs from traditional IDEs, particularly regarding the notebook’s structure and workflow. Jupyter notebooks employ a cell-based execution model that facilitates iterative coding, making them well suited to quickly test hypotheses and iterate on data by reusing existing code cells [73]. A survey among Microsoft data scientists revealed that 94% of the participants considered reusing existing code in Jupyter notebooks to be at least “important” [21]. Users often reuse code through code cloning, where cells are copied within or across notebooks, which can lead to inconsistencies and technical debt [73]. To find reusable code more efficiently, code search tools provide support by helping users locate relevant code cells or workflows [84]. However, proper provenance management is crucial for understanding the origin and evolution of reused code, ensuring that dependencies and execution history are clear [196]. Moreover, reproducibility becomes a challenge in notebooks, as reused code might not produce consistent results if dependencies or execution orders are not properly tracked [177]. In this subsection, we describe these studies that address the challenges and opportunities of reusing code within Jupyter notebooks and provenance.
@@ 5.1.1 Code Cloning
Code cloning in Jupyter notebooks refers to copying and pasting code cells within the same notebook or between different notebooks. Research has shown that code cloning is common in Jupyter notebooks. Over 70% of all notebook code cells on GitHub are exact copies of other cells, and approximately 50% of all notebooks contain no unique cells [63]. These findings are also supported by Yang et al. [198], who noted that about 32% of Jupyter notebooks hosted in GitHub repositories were directly copied from Stack Overflow. The code snippets that are most frequently cloned in Jupyter notebooks mainly relate to visualization (21%), followed by ML (15%) [73]. Among reused code, interproject clones in Jupyter notebooks are far more common than intraproject clones, which is the opposite of the prevalence of code clones in Java code [42]. Furthermore, another study indicated that top-level notebook users (i.e., Grandmasters in Kaggle) are most likely to clone common abstractions such as importing packages, configurations, file IO operations, and showing data [131].
@@ 5.1.2 Reusing Code Snippets by Code Search
Searching for code snippets in Jupyter notebooks can be different from traditional code. While users often seek specific functions or APIs in traditional code, notebook users frequently search for code cells (snippets) that analyze similar data or follow comparable workflows [84]. As highlighted by Li et al. [84], semantic code search is vital in this context, which enables natural language queries that reflect the intended functionality of code cells rather than just API names or keywords. Moreover, the flexible nature of notebooks, where code, data, and results are integrated within a single file, contrasts with the modular organization of traditional code. This integration in notebooks makes it more difficult to isolate functionality and understand the context of individual cells, which imposes challenges for code search such as tracking the logical flow of code and capturing enough context within a code cell [55].
@@
Despite sharing common objectives, code search tools for Jupyter notebooks vary in their methodological approaches. For instance, some tools utilize keyword-based search mechanisms, where users input specific terms related to their queries. This approach, demonstrated in research by Watson et al. [186], allows users to find relevant code by matching their search terms with corresponding code repositories and snippets. On the other hand, certain advanced search tools leverage deep learning techniques to interpret natural language queries. This enables users to search for code snippets more intuitively, using conversational language instead of strict keywords [84]. Similarly, Ragkhitwetsagul et al. [126] proposed Typhon, which integrates traditional information retrieval techniques (BM25) with UniXcoder [49] code embeddings to perform text-to-code matching. In addition, Li and Lv [83] developed a semantic search framework explicitly tailored for Jupyter notebooks, leveraging embeddings generated by large language models (LLMs) from both markdown and code cells.
@@
Additionally, some tools enhance search efficacy by employing graph representations to illustrate code flow relationships within a Jupyter notebook. By mapping the interactions and dependencies between different code cells, these graph-based tools can identify interrelated code cells and suggest connections that the user needs [56]. EDAssistant [85] introduces a methodology for context-aware code search in Jupyter Notebook. This approach prioritizes example notebooks that align with the user’s intent and existing code, assisting users in finding relevant code snippets to reuse and adapt during exploratory data analysis. Similarly focused on cell-level recommendation, Aydin et al. [7] proposed a specialized cell recommendation approach targeting ML notebooks. Furthermore, some search solutions specifically target the data visualization domain by mining, extracting, and cataloging Python functions designed for visualization purposes [10]. This approach leverages reusable visualization snippets from data science notebooks to facilitate access to visualization-related code, thereby promoting the reuse and adaptation of existing visualization workflows.
@@
Good practice #1: Share and explain the data [133]
@@
Good practice #2: Document the process, not just the results [133]
@@
Good practice #3: Record and manage library dependencies [133]
@@
Bad practice #1: Presence of non-executed code cells [116]
@@
Bad practice #2: Out-of-order cell execution [116]
@@
Bad practice #3: Presence of hidden states [115]

### v2

@@ 4.4 Industry Collaborations
Software engineering research on Jupyter notebooks is regularly done in collaboration with industry. Of the 199 primary studies on software engineering research on Jupyter notebooks, 37 (18.8%) were done with industry collaborations. Table 4 presents the industry affiliations of the studies included in our SLR. Microsoft Research has the highest representation among the identified industry collaborations, with a study count of 13. JetBrains and IBM Research have seven and six studies on software engineering research on notebooks, respectively, followed by Microsoft. We analyzed those studies and found that Microsoft Research primarily focused on code management in notebooks head2019managing, subotic2022static, psallidas2022data, mondal2023cell2doc, li2023notable, mcnutt2023design, lin2025interlink. In contrast, IBM Research concentrated on improving code documentation and comprehension liu2021haconvgnn, muller2021data, wang2022documentation, wang2021graph, wang2021makes. JetBrains Research mainly explored tool development aimed at enhancing the execution environment and managing notebook code [titov2024hidden, titov2025observing, grotov2025themisto, titov2022resplit]. Overall, the industry affiliations reflect the widespread interest and impact of software engineering research on Jupyter notebooks in both academia and industry.
@@
Microsoft Research [subotic2022static, psallidas2022data, head2019managing, mondal2023cell2doc, li2023notable, mcnutt2023design, weinman2021fork, chattopadhyay2020whats, chattopadhyay2023make, huang2024code, li2024unlocking, ouyang2024noteplayer, lin2025interlink]
JetBrains Research grotov2022large, titov2022resplit, grotov2024debug, grotov2024untangling, titov2024hidden, titov2025observing, grotov2025themisto
IBM Research [liu2021haconvgnn, muller2021data, wang2021graph, wang2021makes, wang2022documentation, cunha2021context]
Fujitsu Research of America bavishi2021vizsmith, zhu2021restoring, ueno2020ssh
Google Inc [petricek2018wrattler, yin2023natural]
Adobe Research [chen2023whatsnext, carver2025notebookos]
Apple [cheng2024biscuit]
LinkedIn [li2025kernel]
Megagon Labs [kwon2023weedle]
Ploomber [michael2022keeping]
@@ 4.5 Replication Packages
Our systematic literature review discovered a gap in the availability of the resources necessary to replicate the research. Specifically, more than half of the studies (112 out of 199) did not provide publicly accessible source code or installer files of their work. This made it difficult to validate and build on their findings. In contrast, the remaining 87 studies provide source code, installer files, datasets, or a combination thereof to replicate their research results. Upon a closer manual examination of these studies, we found that replication packages for 82 studies are available because they provide the source code with clear instructions to execute. The remaining five studies lacked available replication packages due to the absence of source code or lack of clear execution instructions shankar2022bolt, ritta2022reusing, settewong2022visualize, grotov2025evolving, sparmann2023jugaze. For example, settewong2022visualize shared a GitHub URL [^13] https://github.com/NAIST-SE/VizJupyterNotebooks of their work; however, upon examination, we found neither source code nor installation files. Similarly, no code was found at another GitHub URL [^14] https://github.com/soespa/jugaze [sparmann2023jugaze].
Sharing replication packages of software engineering research on notebooks regularly does not follow open science principles. We found that the dominant platform of choice was GitHub, with 69 studies opting to use it to host their replication packages. Although GitHub is a widely accessible platform, any changes or deletions by repository owners can be a risk to the future accessibility of the replication packages. Best practices for open science in software engineering Mendez_2020 recommend to adopt platforms and repositories designed for the long-term archiving of research outputs. We found that only 17 studies hosted their replication packages on platforms that align with these best practices (i.e., they ensure the permanent archiving of digital artifacts). Of these, 14 were hosted on Zenodo [oli2021automated, ramasamy2023visualising, pimentel2019large, koenzen2020code, quaranta2021kgtorrent, quaranta2022eliciting, aydin2024cellrecommend, golendukhina2024unveiling, mostafavi2024distilkaggle, de2024bug, wang2024why, oden2024integrating, ono2021interactive, grotov2025themisto], two on Figshare [liu2023refactoring, shome2024understanding] and one on OSF Preprints [eckelt2024loops]. The rest has been hosted in a commercial URL [^15] https://www.kave.cc/ [nakamaru2025jupyter].
@@ 4.6 Jupyter Notebook Extensions
We found that a modest 18.1% of the studies (36 of 199) explored the development and use of Jupyter Notebook extensions to enhance the user experience within notebook environments. Despite showing the potential benefits of Jupyter Notebook extensions, 16 studies did not provide available replication packages due to missing links or code cunha2021context, koop2021notebook, muller2021data, watson2019pysnippet, speicher2019notes, verano2022suppose, kwon2023weedle, li2023edassistant, li2023notable, li2022context, fang2025enhancing, jayawardana2024enhancing, kim2024japper, titov2025observing, ueno2020ssh, harden2022exploring. We found that 20 extensions were reusable and integrable into Jupyter notebooks, showcasing a variety of functionalities designed to improve the user experience of Jupyter notebooks wang2022stickyland, bavishi2021vizsmith, kang2021toonnote, pimentel2021understanding, head2019managing, kerzel2023mlprovlab, ramasamy2023visualising, li2023elasticnotebook, perez2024flexible, gadhave2024persist, harrison2024jupyterlab, harden2023there, oden2024integrating, sparmann2023jugaze, cai2025jupyter, nakamaru2025jupyter, tsai2024libyt, eckelt2024loops, rawn2025pagebreaks, husak2023slicito. After replicating and integrating these extensions into Jupyter notebooks, we found that the majority of Jupyter Notebook extensions are designed to enhance the visualization and presentation of notebook cells and their outputs wang2022stickyland, bavishi2021vizsmith, ramasamy2023visualising, kang2021toonnote, head2019managing, perez2024flexible, harden2023there, eckelt2024loops, cai2025jupyter, rawn2025pagebreaks. However, only a few studies tackle more complex challenges related to the core functionalities of Jupyter notebooks, for example, managing data provenance through their extension gadhave2024persist, kerzel2023mlprovlab, allowing live migration of notebook environments li2023elasticnotebook, and notifying potential fairness in data science code harrison2024jupyterlab.
@@
Code reuse and provenance (Section 5.1) Code cloning koenzen2020code, yang2023code, kallen2021jupyter, ritta2022reusing
Reusing code snippets by code search horiuchi2022jupysim, li2021nbsearch, watson2019pysnippet, li2023edassistant, bavishi2021vizsmith, horiuchi2022similarity, aydin2024cellrecommend, ragkhitwetsagul2024typhon, li2024unlocking, li2023dense, bajaj2025tool, li2022context, li2025search
Reproducibility pimentel2021understanding, pimentel2019large, schroder2019reproducible, rule2018ten, wang2020assessing, wang2020restoring, samuel2022computational, ahmad2022reproducible, samuel2021reproducemegit, mietchen2025analyzing, nguyen2025majority
Provenance koop2021notebook, pimentel2015collecting, dao2022runtime, kerzel2023mlprovlab, kerzel2021towards, samuel2018provbook, kery2018interactions, gadhave2024persist, harrison2024jupyterlab, studtmann2023histree, fang2025enhancing, eckelt2024loops, kery2018story
Managing computational environment and workflow (Section 5.2) Empirical studies on workflows lau2020design, raghunandan2023measuring, choetkiertikul2023mining, psallidas2022data, rehman2019towards, koop2017dataflow, subramanian2020casual, golendukhina2024unveiling, zou2024mlpipeline, sparmann2023jugaze
Computational environment in notebooks duan2023jup2kub, cunha2021context, chattopadhyay2020whats, lu2020securing, li2023elasticnotebook, weinman2021fork, ramsingh2024understanding, kinanen2024improving, cao2024jupyter, titov2024hidden, li2024demonstration, zhu2024facilitating, sato2024multiverse, li2024kishu, jayawardana2024enhancing, kim2024japper, zhang2019juneau, tsai2024libyt, christophe2023moon, carver2025notebookos, ueno2020ssh, savira2021writing
Managing dependencies zhu2021restoring, wang2021restoring, fitzpatrick2022davos, wannipurage2022framework
Performance analysis werner2021bridging, werner2024jumper, faenza2024containerized, oden2024integrating, prathanrat2018performance, grotov2025themisto
Readability of notebooks (Section 5.3) Refactoring liu2023refactoring, sato2022comparing, dong2021qualitative, dong2021splitting, titov2022resplit, shankar2022bolt, head2019managing
Non-linear visualization wang2022stickyland, kang2021toonnote, rule2018aiding, chattopadhyay2023make, merino2022making, rawn2025pagebreaks, husak2023slicito, grotov2025evolving, harden2022exploring, harden2023there
Accessibility venkatesh2023notably, nylund2025matplotalt
@@
Documentation of notebooks (Section 5.4) Empirical studies on documentation rule2018design, rule2018exploration, raghunandan2022code, wong2025method
Documentation generation wang2022documentation, wang2021graph, liu2021haconvgnn, wang2021makes, mondal2023cell2doc, muller2021data, li2023notable, lin2023inksight, mostafavi2024beyond, wang2024outlinespark, ouyang2024noteplayer, mostafavi2024can, schmieg2025enhancing, lin2025interlink
Cell header generation venkatesh2023enhancing, venkatesh2021automated, venkatesh2024static, perez2024flexible
@@
Testing and debugging (Section 5.5) Empirical studies on testing and debugging de2024bug, shome2024understanding, wang2024why
Detecting bugs robinson2022error, xin2022provenance, brachmann2020your, patra2022nalin, fangohr2020testing, wang2024using, grotov2024untangling, grotov2024debug, tang2025characterising, jiang2025exploring, titov2025observing
Detecting data leakage yang2022data, subotic2022static, negrini2023static
Visualization in notebooks (Section 5.6) Empirical studies on visualizations agrawal2022understanding, settewong2022visualize, ramasamy2023visualising, wang2023supernova, wootton2024charting, wenskovitch2019albireo, christophe2023moon, tian2025noteflow
Interactive visualization kwon2023weedle, wu2020b2, scully2024design, in2024evaluating, guo2024explainability, cai2025jupyter, ono2021interactive
Best practices (Section 5.7) Following code style standards grotov2022large, wang2020better, siddik2023code, candela2023approach, quaranta2022pynblint, quaranta2024pynblint, speicher2019notes, adams2023comparison, huang2025scientists, cai2025jupyterNotebook
Best practices for collaborative use quaranta2022assessing, quaranta2022eliciting, wang2024dont
Cell execution order (Section 5.8) - oli2021automated, singer2020notes, jiang2022elevating, michael2022keeping, macke2021automating, brown2023facilitating, in2025exploring
AI-based coding assistance for notebooks (Section 5.9) - yin2023natural, brault2023taming, mcnutt2023design, chen2023whatsnext, george2024notebookgpt, huang2024contextualized, cheng2024biscuit, weber2024computational, kimio2024kogi, you2025datawiseagent, fang2025large, jin2025learning, wandel2025pyevalai, jin2025suggesting
Supporting other programming paradigms (Section 5.10) - niephaus2019polyjus, petricek2018wrattler, verano2022suppose, weber2024extending, li2025kernel
Datasets of notebooks (Section 5.11) - grotov2022large, quaranta2022eliciting, li2021nbsearch, liu2021haconvgnn, wang2021restoring, quaranta2021kgtorrent, koenzen2020code, pimentel2019large, raghunandan2022code, de2024bug, liu2023refactoring, ramasamy2023visualising, mostafavi2024distilkaggle, shome2024understanding, golendukhina2024unveiling, huang2024contextualized, grotov2024untangling, wang2024why, aydin2024cellrecommend, ghahfarokhi2024predicting, jin2025suggesting, nakamaru2025jupyter
@@ 5.1 Code Reuse and Provenance
Code reuse in Jupyter notebooks refers to the intentional act of leveraging previously written code cells. Code reuse within Jupyter notebooks differs from traditional IDEs, particularly regarding the notebook’s structure and workflow. Jupyter notebooks employ a cell-based execution model that facilitates iterative coding, making them well suited to quickly test hypotheses and iterate on data by reusing existing code cells koenzen2020code. A survey among Microsoft data scientists revealed that 94% of the participants considered reusing existing code in Jupyter notebooks to be at least “important” chattopadhyay2020whats. Users often reuse code through code cloning, where cells are copied within or across notebooks, which can lead to inconsistencies and technical debt koenzen2020code. To find reusable code more efficiently, code search tools provide support by helping users locate relevant code cells or workflows li2021nbsearch. However, proper provenance management is crucial for understanding the origin and evolution of reused code, ensuring that dependencies and execution history are clear xin2022provenance. Moreover, reproducibility becomes a challenge in notebooks, as reused code might not produce consistent results if dependencies or execution orders are not properly tracked wang2020assessing. In this subsection, we describe these studies that address the challenges and opportunities of reusing code within Jupyter notebooks and provenance.
@@ 5.1.1 Code Cloning
Code cloning in Jupyter notebooks refers to copying and pasting code cells within the same notebook or between different notebooks. Research has shown that code cloning is common in Jupyter notebooks. Over 70% of all notebook code cells on GitHub are exact copies of other cells, and approximately 50% of all notebooks contain no unique cells kallen2021jupyter. These findings are also supported by yang2023code, who noted that about 32% of Jupyter notebooks hosted in GitHub repositories were directly copied from Stack Overflow. The code snippets that are most frequently cloned in Jupyter notebooks mainly relate to visualization (21%), followed by ML (15%) koenzen2020code. Among reused code, interproject clones in Jupyter notebooks are far more common than intraproject clones, which is the opposite of the prevalence of code clones in Java code gharehyazie2019cross. Furthermore, another study indicated that top-level notebook users (i.e., Grandmasters in Kaggle) are most likely to clone common abstractions such as importing packages, configurations, file IO operations, and showing data ritta2022reusing.
@@ 5.1.2 Reusing Code Snippets by Code Search
Searching for code snippets in Jupyter notebooks can be different from traditional code. While users often seek specific functions or APIs in traditional code, notebook users frequently search for code cells (snippets) that analyze similar data or follow comparable workflows [li2021nbsearch]. In this context, semantic code search is vital, which enables natural language queries that reflect the intended functionality of code cells rather than just API names or keywords [li2021nbsearch]. Moreover, the flexible nature of notebooks, where code, data, and results are integrated within a single file, contrasts with the modular organization of traditional code. This integration in notebooks makes it more difficult to isolate functionality and understand the context of individual cells, which imposes challenges for code search such as tracking the logical flow of code and capturing enough context within a code cell [head2019managing].
@@
Despite sharing common objectives, code search tools for Jupyter notebooks vary in their methodological approaches. For instance, some tools utilize keyword-based search mechanisms, where users input specific terms related to their queries. This approach, demonstrated in research by watson2019pysnippet, allows users to find relevant code by matching their search terms with corresponding code repositories and snippets. On the other hand, certain advanced search tools leverage deep learning techniques to interpret natural language queries. This enables users to search for code snippets more intuitively, using conversational language instead of strict keywords [li2021nbsearch]. Similarly, ragkhitwetsagul2024typhon proposed Typhon, which integrates traditional information retrieval techniques (BM25) with UniXcoder [guo2022unixcoder] code embeddings to perform text-to-code matching. In addition, li2024unlocking developed a semantic search framework explicitly tailored for Jupyter notebooks, leveraging embeddings generated by large language models (LLMs) from both markdown and code cells.
@@
Additionally, some tools enhance search efficacy by employing graph representations to illustrate code flow relationships within a Jupyter notebook. By mapping the interactions and dependencies between different code cells, these graph-based tools can identify interrelated code cells and suggest connections that the user needs [horiuchi2022jupysim]. EDAssistant [li2023edassistant] introduces a methodology for context-aware code search in Jupyter Notebook. This approach prioritizes example notebooks that align with the user’s intent and existing code, assisting users in finding relevant code snippets to reuse and adapt during exploratory data analysis. Similarly focused on cell-level recommendation, aydin2024cellrecommend proposed a specialized cell recommendation approach targeting ML notebooks. Furthermore, some search solutions specifically target the data visualization domain by mining, extracting, and cataloging Python functions designed for visualization purposes [bavishi2021vizsmith]. This approach leverages reusable visualization snippets from data science notebooks to facilitate access to visualization-related code, thereby promoting the reuse and adaptation of existing visualization workflows.
Continuing this trend, new approaches have emerged to improve code search and reuse in notebooks. For instance, a system called DeCNR was developed to fuse sparse and dense retrieval models to effectively locate relevant computational notebooks [li2023dense]. In addition, the authors introduced a dedicated evaluation dataset of notebook queries with relevance judgments to enable fair benchmarking of search techniques. Similarly, a context-aware notebook search framework was developed to help researchers seamlessly discover external notebooks relevant to their current analysis, integrating semantic search capabilities into a Jupyter-based virtual research environment [li2022context]. In educational settings, bajaj2025tool presented JBEval, a tool that identifies similar code and text blocks across a collection of notebooks to detect reuse or plagiarism, providing similarity scores for code and text and visualizing identical segments for further inspection. Furthermore, a unified search system was introduced that allows users to query multiple types of research assets (including code snippets, datasets, and documentation) directly from within a notebook, reflecting a move toward richer, context-integrated search functionalities in Jupyter environments [li2025search].
@@
Good practice #1: Share and explain the data [rule2018ten]
@@
Good practice #2: Document the process, not just the results [rule2018ten]
@@
Good practice #3: Record and manage library dependencies [rule2018ten]
@@
Bad practice #1: Presence of non-executed code cells [pimentel2021understanding]
@@
Bad practice #2: Out-of-order cell execution [pimentel2021understanding]
@@
Bad practice #3: Presence of hidden states [pimentel2019large]